OpenStudy (m.d.saxon):

58a +28 -15 =651 95a +354 -25 +3a =15147 What is the value of a? @Elsa213 @this-emo-kid9 @♥THIS_EMO_GIRL♥ @RyGuy @emmigrace222 @emjaye2k99 @goformit100

OpenStudy (anonymous):

58a = 651-15+28 so 58a = 664 so a= 664/58 = 11.4 98a = 15147 - 354 + 25 so a = 14818/98 = 151.2 are they both supposed to be the same? because they come out different. or are they two seperate questions?

OpenStudy (m.d.saxon):

nope ur wrong

OpenStudy (anonymous):

is a for both questions supposed to be the same?

OpenStudy (anonymous):

if so, could you just quickly check if maybe you mistyped a number ?

OpenStudy (m.d.saxon):

I fixed it

OpenStudy (anonymous):

are both a's supposed to be equal?

OpenStudy (m.d.saxon):

yes